OpenStudy (anonymous):

Write an equation in y = mx + b form for the table. gallons of gas used 1 2 3 4 miles driven 26 52 78 104 Answer y = 26x y = 26x + 26 y = 26x + 1 y = 26x -1

OpenStudy (anonymous):

y-26=m(x-1) y=26(x-1)+26 y=26x-26+26 y=26x